OpenStudy (anonymous):

Element L has an average atomic mass of 59.9994 amu and is composed of two isotopes. The first has a mass of 61.2012 amu and makes up 28.970% of a standard sample. What is the mass of the second isotope?

OpenStudy (anonymous):

I need help please

OpenStudy (australopithecus):

Hint use the formula for Average to come up with the first equation Instructions: \[Average = \frac{a+ b}{n}\] where, n = the number of elements being averaged For example the average of 4, 4, 4, 4, 4, there are 5 elements, so (4 + 4 + 4 + 4 + 4)/5 = 4 the average of 4, 4, 4, 4, 4 elements, so (4 + 4 + 4 + 4)/4 = 4 In your problem you are given the average molecular mass, told there are two elements and you are also told one of the molecular masses of the elements

OpenStudy (anonymous):

please which one is the answer

OpenStudy (australopithecus):

Can you solve for x? This isnt even a system of linear equations problem 59.9994 = (61.2012 + x)/2

OpenStudy (abmon98):

59.9994=(61.2012* 28.970/100)+(x*(100- 28.970)/100) 59.9994=17.7299+0.7103x x=59.9994-17.7299/0.7103 x=59.509